Compare the reactor volume required to achieve 95% contaminant removal efficiency for a CMFR vs. a PFR for the following cond It's not 840, 5320, or 6.34
Answer #1

According to the given data following is the answer to the problem. If it isn't what required there will be something wrong with the given data.

Assumptions:

  1. Steady State condition
  2. First order reaction
  3. No outside gain or loss due to percolation or evaporation
  4. Non conservative compound

Given:

Q = 14 m3/day

k = 0.05 /day

n = 95%

Answer:

n = [ 1- ( C0/Cin) ] * 100

= 95 %

Therefore,

C0/Cin = 1 - (95/100)

= 0.05

For CMFR,

C0/Cin = [1/ ( 1+K*a )]

0.05 = [ 1/ ( 1+ 0.05 * a) ]

a = 380 days

VCMFR = a * Q

= 380 days * 14 m3/day

= 5320 m3

For PFR,

C0/Cin = e - (ka)

0.05 = e - 0.05 * a

    a = 59.915 days

VPFR = a * Q

   = 59.915 days * 14 m3/day

= 838.81 m3

Dimensionless Ratio,

VCMFR / VPFR = 5320 / 838.81

= 6.342

Hence the required ratio is 6.342